# Nicholas Michel    Mohamed E. Najd    Ghada Almashaqbeh

### v3

@@ Abstract
Automated market makers (AMMs) are a form of decentralized cryptocurrency exchanges that have attracted huge interest lately. They are considered a prime example of Decentralized Finance (DeFi) applications, a large category under Web 3.0. Their popularity and high trading activity have resulted in millions of on-chain transactions leading to serious scalability issues in terms of throughput and on-chain state size. Existing scalability solutions, when employed in the context of AMMs, are either ineffective due to their large overhead, or suffer from security and centralization issues.
@@
In this paper, we address these challenges by utilizing a new sidechain architecture as a layer 2 solution, building a system called $\mathsf{ammBoost}$. Our system reduces the amount of on-chain transactions, boosts throughput, and supports blockchain pruning. We devise several techniques to enable layer 2 processing while preserving the correct and secure operation of AMMs. These include a functionality-split and layer 2 traffic summarization paradigm, an epoch-based deposit mechanism, and pool snapshot-based and delayed token-payout trading. We also build a proof-of-concept of $\mathsf{ammBoost}$ for a Uniswap-inspired use case to empirically evaluate performance. Our experiments show that $\mathsf{ammBoost}$ decreases the gas cost by 96.05% and the chain growth by at least 93.42%, and that it can support up to 500x of the daily traffic volume observed for Uniswap in practice.
1 Introduction
@@
Cryptocurrencies and blockchain technology provide an innovative model that led to new applications and research frontiers, as well as reshaping the Internet and its digital services (under what is called Web 3.0). Decentralized Finance (DeFi) is a large category under Web 3.0 in which blockchains are used to transform traditional financial services, which are usually centrally managed, into fully decentralized ones. Many of these systems operate in an open-access model, thus removing market entrance barriers for customers, and enabling a transparent and intermediary-free interaction. Smart contracts strengthen this model by providing an automated way to negotiate contract terms and enforce agreements.
@@
Automated market makers (AMMs) are considered a prime example of DeFi services [71]. They build a platform for automated token trading by establishing liquidity pools for token pairs. An AMM is implemented as a decentralized application (dApp); a set of smart contracts on top of a smart contract-enabled blockchain—where Ethereum is the dominant choice so far, that support operations for trading and liquidity management, such as swaps, mints, burns, and collects. AMMs are a huge industry with a total monthly trading volume of \$46 - \$95 billion (during the first half of 2023), and an estimated total market cap of nearly \$16 billion as of January 2024 [24]. Many popular AMMs are deployed in practice and widely used. Examples include Uniswap [27], Curve [8], DODO [9], and Sushiswap [22], which during the first half of 2023 commanded around 62-71%, 6.38-14.01%, 4.04-7.16%, and 0.69-3.67% of the top ten AMMs market share, respectively [1].
@@
Challenges. At the same time, AMMs are a huge scalability problem. The popularity and high trading activity of AMMs led to serious efficiency problems since they produce a massive number of on-chain transactions. On the one hand, this increases the underlying blockchain size or storage overhead, and on the second hand, it incurs large (gas) fees. This large workload does not only amplify state storage cost, but also transaction processing/confirmation delays due to the low throughput of blockchains (Ethereum’s throughput is around 12 transaction/sec on average [11]). [^1] A dApp on-chain storage includes the state of the smart contract account of this dApp, and the transaction history recorded on the blockchain that produced this state. A misconception about on-chain storage cost is counting only the contract latest state while ignoring the permanent on-chain storage of these transactions.
@@
Concretely, based on the traffic analysis that we conducted for 2023 (see Appendix 0.C), Uniswap V3 users produced 20 million transactions on Ethereum in 2023 (transaction sizes range from $400$ bytes to $3000$ bytes). This translates to adding around $20.2$ GB to the Ethereum blockchain. In its year of deployment (Nov 2018), Uniswap V1 generated 34,000 transactions, and in 2023, Uniswap’s various versions on Ethereum generated around 80 million transactions, leading to 231,7% increase in transaction volume. These numbers indicate that the scalability problem of AMMs is amplified over the years. This does not only impact the AMM itself, but also other dApps deployed on the underlying blockchain. Such contention drives users to put in high transaction fees so that miners would prioritize their transactions.
@@
The challenge is to handle this scalability problem without impacting security (by, e.g., employing trusted third parties or weakening the security of consensus, which introduces new threat vectors).
@@
Limitations of prior work. Improving blockchain scalability is an active research area. Many solutions have been proposed; some of them target layer 1 improving on the consensus itself, such as sharding [42, 73, 53], while others target layer 2 allowing for some form of off-chain processing, such as payment channels and networks [43] and rollups [16, 31]. However, when it comes to AMMs, applying these solutions impacts performance and security, and may not even cut the storage cost. In sharding, localized workload division policies are used to reduce cross-shard transactions. For smart contract-enabled blockchains, this means that a dApp (so the whole AMM) will be contained in one shard [33, 66], thus parallel processing among shards is not utilized. Others [60] use static analysis to shard dApps by splitting them into commuting functional units that can be executed in any order, and assigning each unit to a shard. This cannot work for AMMs since their (per pool) operation is sequential. Distributing liquidity pools among shards has been proposed in [62], but the reliance on locked cross-shard transactions (to support multi-swaps and arbitraging) may degrade the AMM performance. All these sharding solutions log all transactions on-chain (i.e., the shards), so they do not cut the storage cost. [^2] Ledger pruning for sharding has been proposed in [53, 67] but in the UTXO model, i.e., as in Bitcoin; it is not for the account model (as in Ethereum) that represents smart contract-enabled blockchains on top of which AMMs are deployed.
@@
Layer 2 solutions that allow computations, i.e., beyond just currency transfer as in payment networks [43], have limitations. Optimistic rollups have long contestation periods that may reach one week as in Optimism and Arbitrum [16, 49]. Thus, a user cannot act based on the submitted state changes immediately, but has to wait until the end of the contestation period to ensure that the submitted results are valid. Moreover, they have security issues—verifiers (who validate the submitted changes during the contestation period) could be centralized trusted entities [65], while incentive compatibility of non-trusted verifiers is still an open question [56, 5, 17] which may lead to adopting incorrect ledger state changes. Zero-knowledge (ZK)-rollups [21, 39, 38] are costly; proof generation may take several minutes and it becomes worse when attesting to complex transactions. They also may have a long transaction confirmation delay that may reach 24 hours as in zkSync Era [31]. This impacts transaction processing and confirmation delays, forcing the users to wait longer for their transactions to be finalized. Not to mention that many of the used ZK systems require a trusted setup. [^3] Another instance of layer 2 solutions resorted to employing a centralized settlement party that matches trades and generate ZK proofs to prove settlement correctness [50]. Thus, this solution is a form of ZK-rollups that employs a single settlement party.
@@
Sidechains [36, 46, 47, 51] are another type of layer 2 solutions that can improve scalability. Despite their potential, existing efforts mostly focus on two-way peg, i.e., currency transfer between the sidechain and the mainchain, and all of them are considered independent sidechains. That is, each chain has its own transactions, miners, and tokens. Such independence and the focus on two-way peg limit the performance gains that can be achieved, and do not allow for workload sharing between the chains. Moreover, none of these solutions allow pruning stale records. [^4] In parallel to these academic efforts, many industrial initiatives have explored the utility of sidechains [61, 6, 70, 2, 55, 30]. Most of them focused on independent sidechains and two-way peg. Cosmos [6] allows some form of data exchange via publishing events and commitments on the destination chain. This worsens the storage problem due to data duplicates, and there is no support for blockchain pruning. Other instances in practice resorted to operating the whole AMM on an independent sidechain, i.e., as a separate system from the underlying smart contract-enabled blockchain. For example, Polygon [18] operates a fast EVM-compatible sidechain running along with Ethereum. The whole AMM is run on this sidechain and tokens can be transferred to Ethereum using bridges or atomic swap techniques. Isolating an AMM on an independent sidechain impacts composability with other dApps running on the mainchain (which is Ethereum in this case), limits interaction with the mainchain to merely currency transfers, and complicates system design and its security due to the involvement of bridges. Not to mention that this solution just moves the on-chain storage cost to the sidechain, which will have scalability issues on its own as the AMM user population grows.
@@
A new approach. Sidechains seem to have (so far under-utilized) potential in building an effective layer 2 solution to promote scalability. This has been observed in [57], who proposed a framework called chainBoost with a new sidechain architecture that has a mutual-dependence relation with the mainchain, thus permitting workload sharing and arbitrary data exchange, as well as blockchain pruning. chainBoost targets resource markers—Web 3.0 systems that offer decentralized digital services, e.g., Filecoin [12] and Livepeer [15]. chainBoost directs all heavy/frequent service-related traffic to the sidechain, which in turn processes this traffic and produces concise summaries of the state changes that are used to sync the mainchain. Once these summaries are confirmed on the mainchain, the temporary blocks containing the actual transactions on the sidechain are pruned. The empirical results in [57] show substantial performance gains in terms of blockchain size, transaction confirmation delays, and throughput. All of these are achieved without compromising security and while keeping the mainchain as the single truth of the system state.
@@
These advantages motivated us to explore the following: Can we control the state growth of AMMs, and boost their throughput, in a secure and low-overhead way, and without isolating the AMM on a separate blockchain, using dependent-sidechains?
1.1 Our Contributions
@@
We answer this question in the affirmative and propose $\mathsf{ammBoost}$; a secure storage control and throughput boosting solution for AMMs. In particular, we make the following contributions.
@@
System design. $\mathsf{ammBoost}$ introduces a novel approach for dividing the AMM functionality into two modules: one that resides on the mainchain (i.e., the underlying smart contract-enabled blockchain) and another that is operated by the sidechain. In particular, $\mathsf{ammBoost}$ offloads processing most transactions (swaps, mints, collects, and burns) to the sidechain, and minimizes the functionality remaining on the mainchain. The latter is encapsulated in a base smart contract called $\mathsf{TokenBank}$, that manages the actual tokens by tracking only the transaction summaries produced by the sidechain. It also includes all operations that must happen in real-time on the mainchain, such as flash loans.
@@
$\mathsf{ammBoost}$ solves several challenges related to applying dependent sidechains to AMMs. The chainBoost framework assumes a mutual-dependency relation between the main and side chains; both operate in the same domain and have the same transaction format, services, and miner population. This is not the case for $\mathsf{ammBoost}$; the AMM is merely a dApp deployed at the application layer, so it does not modify how the mainchain protocol works. Thus, the mainchain miners do not maintain the sidechain as in chainBoost, even they might not be aware of its existence. However, $\mathsf{ammBoost}$’s sidechain is impacted by the mainchain since the tokens and the AMM state (and some core functionalities) are on the mainchain. This means means that $\mathsf{ammBoost}$ introduces a unidirectional dependency relation: the sidechain is impacted by the mainchain but not vice versa. Furthermore, the sidechain should process all trading activities without a custody of the actual tokens, and must ensure that only transactions for which issuing users own tokens on the mainchain are accepted.
@@
We resolve these challenges by introducing several techniques. First, we require the AMM to have its own miners to maintain the sidechain. So, like any blockchain, these miners are assumed to have a mining power with an honest majority, they need to build a Sybil-resistant identity by, e.g., using a proof-of-stake approach, and they will be rewarded for maintaining the sidechain using, e.g., the AMM native token. Second, we introduce epoch-based deposits, where a user has to deposit on the mainchain the anticipated amount of tokens needed to back up her activities (or issued transactions) during an epoch on the sidechain. Third, we introduce pool snapshot-based and delayed token-payout trading. That is, the pool token balances are retrieved from the mainchain at the beginning of the epoch, which are used to compute trade prices processed on the sidechain. These balances evolve during the epoch based on the processed transactions processed. Users can use newly accrued tokens in trading since all balances are tracked, but they cannot withdraw the actual tokens directly. This is because the sidechain does not hold actual tokens; token payouts and deposit leftover refunds happen at the end of the epoch when $\mathsf{TokenBank}$ is synced.
@@
Security analysis. We analyze the security of $\mathsf{ammBoost}$ showing that it preserves the correct and secure operation of the AMM.
@@
Implementation and evaluation. We also build a proof-of-concept implementation for a Uniswap-inspired use case, and conduct experiments to empirically evaluate the performance gains, in terms of blockchain size, confirmation delay, and throughput, that $\mathsf{ammBoost}$ can achieve. Our experiments show that $\mathsf{ammBoost}$ achieves a 96.05% gas cost reduction and 93.42% chain growth reduction (when compared to a Uniswap version deployed on the Sepolia testnet). Our experiments demonstrate that $\mathsf{ammBoost}$ can support large traffic volumes, on the order of up to 500x of Uniswap’s daily transaction volume. We also study impact of various configuration parameters on the performance gains of $\mathsf{ammBoost}$.
@@
Although the focus of this work is on controlling the storage overhead and boosting throughput of AMMs, we believe that $\mathsf{ammBoost}$’s paradigm can enable more optimizations for AMMs, e.g., integration of privacy-preserving techniques. Also, $\mathsf{ammBoost}$ could be beneficial for other DeFi applications, and dApps in general, as it introduces a framework for operating application-specific sidechains interacting with smart contract-enabled blockchains. We leave exploring such directions as part of our future work.
2 Background

### v4

# ammBoost: State Growth Control for AMMs

@@ Abstract
Automated market makers (AMMs) are a prime example of Web 3.0 applications. Their popularity and high trading activity led to serious scalability issues in terms of throughput and state size. In this paper, we address these challenges by utilizing a new sidechain architecture, building a system called $\mathsf{ammBoost}$. $\mathsf{ammBoost}$ reduces the amount of on-chain transactions, boosts throughput, and supports blockchain pruning. We devise several techniques to enable layer 2 processing for AMMs, including a functionality-split and layer 2 traffic summarization paradigm, an epoch-based deposit mechanism, and pool snapshot-based and delayed token-payout trading. We also build a proof-of-concept for a Uniswap-inspired use case to empirically evaluate performance. Our experiments show that $\mathsf{ammBoost}$ decreases the gas cost by 96.05% and the chain growth by at least 93.42%, and that it can support up to 500x of the daily traffic volume of Uniswap. We also compare $\mathsf{ammBoost}$ to an Optimism-inspired solution showing a 99.94% reduction in transaction finality.
Index Terms:
AMMs, sidechains, scalability, layer-2 solution.
I Introduction
@@
Cryptocurrencies and blockchain technology provide an innovative model that ignited a movement towards the next generation of the Internet—Web 3.0. Decentralized Finance (DeFi) is a large category under Web 3.0 in which blockchains are utilized to transform traditional financial services, which are usually centrally managed, into fully decentralized ones. Many of these systems operate in an open-access model, thus removing entrance barriers for users, and enabling a transparent and intermediary-free interaction.
@@
Automated market makers (AMMs) are considered a prime example of DeFi services [1]. They build a platform for automated token trading by establishing liquidity pools for token pairs. An AMM is implemented as a decentralized application (dApp) composed of a set of smart contracts deployed on a smart contract-enabled blockchain—where Ethereum is the dominant choice so far. It supports operations to handle trading and liquidity management, such as swaps, mints, burns, and collects. AMMs are a huge industry with a total monthly trading volume of \$46--\$95 billion (during the first half of 2023), and an estimated total market cap of \$16 billion as of January 2024 [2]. Many popular AMMs are deployed in practice and widely used. Examples include Uniswap [3], Curve [4], DODO [5], and Sushiswap [6], which command a large portion of the AMM market share [7].
@@
Challenges. At the same time, AMMs are a huge scalability problem. Their popularity and high trading activity led to serious efficiency problems since AMMs produce a massive number of on-chain transactions. On the one hand, this increases the underlying blockchain size, and on the second hand, it incurs large (gas) fees. This large workload does not only amplify state storage cost, [^1] A dApp on-chain storage includes the state of the dApp’s smart contract, and the transaction history recorded on the blockchain that produced this state. A misconception about on-chain storage is counting only the contract latest state while ignoring the permanent on-chain storage of these transactions. but also transaction processing/confirmation delays due to the low throughput of blockchains (on Ethereum, average throughput is around 12 transaction/sec [8], and 62% and 23% of all transactions have a delay of at least 5 min and 30 min, respectively [9]).
@@
Concretely, based on the traffic analysis that we conducted for 2023 (see Appendix D), Uniswap V3 produced 20 million transactions on Ethereum in 2023 (transaction sizes range from $400$ bytes to $3000$ bytes). This translates to adding around $20.2$ GB to the Ethereum blockchain. In its year of deployment (Nov 2018), Uniswap V1 generated 34,000 transactions, and in 2023, Uniswap’s various versions on Ethereum generated around 80 million transactions, leading to 2,353% increase in transaction volume. These numbers indicate that the scalability problem of AMMs is amplified over the years. This does not only impact the AMM itself, but also other dApps deployed on the underlying blockchain. Such contention drives users to put in high transaction fees so that miners would prioritize their transactions.
@@
Thus, the challenge is how to address the scalability problem of AMMs without impacting security (e.g., without employing trusted third parties or weakening consensus security).
Limitations of prior work. Existing blockchain scalability solutions target layer 1, such as sharding [10, 11, 12, 13], or layer 2, such as rollups [14, 15], that enable off-chain processing. Applying these solutions to AMMs impacts performance and security, and does not even cut storage costs.
@@
In sharding, localized workload division is used to reduce cross-shard transactions. For smart contract-enabled blockchains, this means that a dApp (so the whole AMM) will reside on one shard [16, 17], thus parallel processing among shards is not utilized. Others [18] use static analysis to shard dApps by splitting them into commuting functional units that can be executed in any order, and assigning each unit to a shard. This cannot work for AMMs since their (per pool) operation is sequential. Distributing liquidity pools among shards [19] relies on locked cross-shard transactions (to support multi-swaps and arbitraging), which may degrade the AMM performance.
Solution* Type Throughput Token Payout Delay Liquidity Withdrawal Overhead Decentralized Mainchain Storage
Uniswap Optimism [20, 21] Optimistic Rollup 0.6 tx/s 7 days 4 tx (including Burn) No Batch-txn transcript
Unichain [22, 23] Optimistic Rollup 1.92 tx/s 7 days 4 tx (including Burn) Yes Batch-txn transcript
ZKSwap [24, 25, 26, 27] ZK-rollup 8 - 25 tx/s 3-24 hrs 2-3 tx (including Burn) No State changes
$\mathsf{ammBoost}$ Sidechain 138.06 tx/s 346.49 s 1 (Burn) tx Yes State changes
    For Unichain, average throughput reported for its deployment [28]. ZKSwap is deployed on top of ZKSync-Era [15], so reported metrics are for the latter [29, 26].
TABLE I: Comparison between $\mathsf{ammBoost}$ and rollup solutions.
@@
Layer 2 solutions that allow computations, i.e., beyond just currency transfer as in payment networks [30], also have limitations. Optimistic rollups have long contestation periods that may reach one week as in Optimism and Arbitrum [14, 31]. Thus, a user has to wait until the end of this period to ensure that submitted results are valid, and withdraw tokens accrued from trades [20, 32]. Moreover, they have security issues; verifiers (who validate submitted results during contestation) could be centralized trusted entities [33], while decentralized verifier networks have incentive compatibility issues, i.e., configuring a proper incentive model that encourage verifies to vet the results is still an open question [34, 35, 36], which may lead to adopting incorrect ledger state changes. Also, optimistic rollups are vulnerable to fraud proof denial attacks causing dispute transactions to be delayed, and then rejected, as the contestation period is over [37].
Zero-knowledge (ZK)-rollups [38, 39, 40] are costly; proof generation is computationally heavy (a prover in zksync era can cost 1365 USD per month [41]) and it becomes worse when attesting to complex transactions, and many require a trusted setup. They also may have a long transaction confirmation delay that may reach up to 24 hours as in zkSync Era [15] (with a recent upgrade cutting that to 3 hrs [25]). Such delays force users to wait longer for their transactions to be finalized, which negatively impact the operation of AMMs. Some solutions employ a centralized settlement party for trade matching with ZK proofs to prove settlement correctness [42], so they are a form of centralized ZK-rollups.
@@
Another instance of layer 2 solutions is sidechains [43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Existing efforts mostly focus on two-way peg, i.e., currency transfer between the chains, and only consider independent sidechains. That is, each chain has its own transactions, miners, and tokens. This narrow focus limits the performance gains that sidechains can achieve, and prevents workload sharing between the chains. Cosmos [48] allows some form of data exchange using data duplicates, thus amplifying storage costs. For AMMs, Polygon [53] operates an EVM-compatible sidechain hosting the whole AMM, so it is separated from the mainchain, where tokens can be transferred to Ethereum using bridges. Isolating an AMM on a separate blockchain impacts composability with other dApps, and limits interaction with the mainchain to merely currency transfers. Also, it just moves the on-chain storage cost to the sidechain, which will have scalability issues on its own as the AMM user population grows.
In terms of blockchain pruning, pruning for sharding [13, 54] targets the UTXO model; it is not for the account model underlying smart contract-enabled blockchains. Also, [13] moves the responsibility of storing transactions (to prove their existence) to the clients. Snapshot-based pruning solution [55, 56, 57, 58] either target the UTXO model [57, 56], support only permissioned blockchains [55], or require archival nodes that store the full blockchain as in Bitcoin and Ethereum [59, 60, 58]. Optimistic rollups do not prune since the processed transaction batch is stored on-chain [61, 62], or on a third-party data availability layer [63, 64], to allow verifiers to verify the submitted results [41, 61]. ZK-rollups either store the batch on-chain [65] or avoid that at the expense of more expensive ZK proofs [41, 66]. For sidechains, none of the existing solutions support pruning.
We present a comparison with different layer 2 scalability solutions in Table I.
@@
A new approach. Existing solutions around sidechains, due to their narrow focus, fall short in realizing the full potential of sidechains in building an effective layer 2 scalability solution. This has been observed in [67], who proposed a framework called chainBoost with a new sidechain architecture that has a mutual-dependence relation with the mainchain, thus permitting workload sharing, arbitrary data exchange, and blockchain pruning. chainBoost targets resource markets—Web 3.0 systems that offer decentralized digital services [68, 69]. It directs all heavy/frequent service-related traffic to the sidechain, which in turn processes this traffic and produces summaries of the state changes that are used to sync the mainchain. Once these summaries are confirmed on the mainchain, the temporary blocks containing the actual transactions on the sidechain are pruned. The empirical results in [67] show substantial performance gains in terms of blockchain size, confirmation delays, and throughput. All of these are achieved without compromising security and while keeping the mainchain as the single truth of the system state.
@@
These advantages motivated us to ask whether we can utilize dependent-sidechains to control the state growth of AMMs, and boost their throughput, in a secure and low-overhead way without isolating the AMM on a separate blockchain.
I-A Our Contributions
@@
We answer this question in the affirmative and propose $\mathsf{ammBoost}$; a secure storage control and throughput boosting solution for AMMs. We make the following contributions.
@@
System design. $\mathsf{ammBoost}$ introduces a novel approach for dividing the AMM functionality into two modules: one that resides on the mainchain and another that is operated by the sidechain. In particular, $\mathsf{ammBoost}$ offloads processing most transactions (swaps, mints, collects, and burns) to the sidechain, and minimizes the functionality remaining on the mainchain. The latter is encapsulated in a base smart contract, called $\mathsf{TokenBank}$, that manages the actual tokens by tracking only the transaction summaries produced by the sidechain. It also includes all operations that must happen in real-time on the mainchain, such as flash loans.
@@
$\mathsf{ammBoost}$ solves several challenges related to combining dependent sidechains with AMMs. The chainBoost framework assumes a mutual-dependency relation between the main and side chains; both are aware of each other and operate in the same domain, i.e., same transaction format, protocol, and miners. This is not the case for $\mathsf{ammBoost}$; the AMM is merely a dApp deployed at the application layer, so applying chainBoost as is requires modifying the mainchain protocol which leads to hard forks. However, $\mathsf{ammBoost}$’s sidechain is impacted by the mainchain since the tokens and the AMM state are on the mainchain. This means that $\mathsf{ammBoost}$ exhibits a unidirectional dependency relation. Furthermore, the sidechain does not hold custody of actual tokens, these reside on the mainchain, but it has to process transactions correctly and accept only these for which issuing users own tokens on the mainchain. Moreover, since the mainchain miners in $\mathsf{ammBoost}$ do not track the sidechain, sync-transaction authentication is needed so that only syncs issued by the rightful epoch committee in $\mathsf{ammBoost}$ are accepted by $\mathsf{TokenBank}$.
@@
We resolve these challenges by introducing several techniques. First, we require the AMM to have its own miners to maintain the sidechain. So, like any secure blockchain, these miners have a mining power (with honest majority), and they will be rewarded for maintaining the sidechain using, e.g., AMM’s native token. This allows enjoying the scalability benefits of $\mathsf{ammBoost}$ without introducing core changes to the mainchain protocol that lead to hard forks. Second, to ensure that users owns actual tokens for trading, we introduce epoch-based deposits, where a user deposits on the mainchain the anticipated amount of tokens needed to back up her issued transactions during an epoch on the sidechain.
Third, to allows trading on the sidechain without token custody, we introduce pool snapshot-based and delayed token-payout trading. That is, the pool token balances are retrieved from the mainchain at the beginning of the epoch, which are used to compute trade prices processed on the sidechain. These balances evolve during the epoch based on the transactions processed. Users can directly use newly accrued tokens in trading since all balances are tracked, but they cannot withdraw actual tokens until the epoch ends. This is because the sidechain does not hold actual tokens; token payouts and deposit leftover refunds happen when $\mathsf{TokenBank}$ is synced. Fourth, to ensure that only legit sync-transactions accepted on the mainchain, we introduce a sync authentication mechanism that combines quorum certificates and threshold signatures ensuring that only syncs issued by the rightful sidechain committees are accepted by $\mathsf{TokenBank}$.
A large advantage of $\mathsf{ammBoost}$’s approach to sidechains is maintaining dApp composability. That is, the base smart contract on the mainchain maintains the application state. Thus, interactions with other dApps happen on the mainchain, e.g., as in the original Ethereum setting.
@@
Security analysis. We analyze the security of $\mathsf{ammBoost}$ showing that it is a secure scalability solution as it preserves the correct and secure operation of the underlying AMM.
@@
Implementation and evaluation. We also build a proof-of-concept implementation for a Uniswap-inspired use case, and conduct experiments to empirically evaluate the performance gains, in terms of blockchain size, confirmation delay, and throughput, that $\mathsf{ammBoost}$ can achieve. Our experiments show that $\mathsf{ammBoost}$ achieves a 96.05% gas cost reduction and 93.42% chain growth reduction (when compared to a Uniswap version deployed on the Sepolia testnet). Our experiments demonstrate that $\mathsf{ammBoost}$ can support large traffic volumes, on the order of up to 500x of Uniswap’s daily transaction volume. We also study impact of various configuration parameters on the performance gains of $\mathsf{ammBoost}$, and compare it to an Optimism-inspired solution showing a 99.94% reduction in transaction finality.
@@
Lastly, we believe that $\mathsf{ammBoost}$’s paradigm can enable more optimizations for AMMs, e.g., integration of privacy-preserving techniques. Also, $\mathsf{ammBoost}$ could be beneficial for other DeFi applications, and dApps in general, as it introduces a framework for operating application-specific sidechains interacting with smart contract-enabled blockchains. We leave exploring such directions as part of our future work.
II Background and Challenges